# PAC-Bayesian Reinforcement Learning Trains Generalizable Policies

### v1

@@ B.6 A note on the Markov assumption for Bellman errors
Appendix C REINFORCE Trick for Policy-Level Gradients
@@
Theorem C.1 (Policy-Level REINFORCE).
For a posterior distribution $\rho=\mathcal{N}(\upsilon,\operatorname{diag}(\sigma^{2}))$ over policy parameters $\theta$, the gradient of the expected return can be computed as:
$\nabla_{\upsilon,\sigma}\mathbb{E}_{\theta\sim\rho}\left[\mathbb{E}[R_{\tau}|\pi_{\theta}]\right]=\mathbb{E}_{\theta\sim\rho}\left[\nabla_{\upsilon,\sigma}\log\mathbb{P}_{\upsilon,\sigma}(\theta)\cdot\mathbb{E}[R_{\tau}|\pi_{\theta}]\right]$ (37)
@@
Starting with the gradient of the expected return of a policy over the posterior distribution:
$\nabla_{\upsilon,\sigma}\mathbb{E}_{\theta\sim\rho}\left[\mathbb{E}[R_{\tau}|\pi_{\theta}]\right]$ $=\nabla_{\upsilon,\sigma}\sum_{\theta_{i}}\mathbb{P}_{\upsilon,\sigma}(\theta_{i})\sum_{a}\pi_{\theta_{i}}(a)\underset{q(a)}{\underbrace{\mathbb{E}[R_{\tau}|A_{\tau}=a]}}$ (38)
Moving the gradient inside the summation:
$=\sum_{\theta_{i}}\nabla_{\upsilon,\sigma}\mathbb{P}_{\upsilon,\sigma}(\theta_{i})\sum_{a}\pi_{\theta_{i}}(a)q(a)$ (39)
where $q(a)=\mathbb{E}[R_{\tau}|A_{\tau}=a]$ is the state-action value for notational convenience.
Multiplying by $\frac{\mathbb{P}_{\upsilon,\sigma}(\theta_{i})}{\mathbb{P}_{\upsilon,\sigma}(\theta_{i})}$:
$=\sum_{\theta_{i}}\left[\left(\sum_{a}\pi_{\theta_{i}}(a)q(a)\right)\frac{\mathbb{P}_{\upsilon,\sigma}(\theta_{i})}{\mathbb{P}_{\upsilon,\sigma}(\theta_{i})}\nabla_{\upsilon,\sigma}\mathbb{P}_{\upsilon,\sigma}(\theta_{i})\right]$ (40)
Applying the log-derivative trick $\nabla_{\upsilon,\sigma}\mathbb{P}_{\upsilon,\sigma}(\theta_{i})=\mathbb{P}_{\upsilon,\sigma}(\theta_{i})\nabla_{\upsilon,\sigma}\log\mathbb{P}_{\upsilon,\sigma}(\theta_{i})$:
$=\sum_{\theta_{i}}\left[\mathbb{P}_{\upsilon,\sigma}(\theta_{i})\left(\sum_{a}\pi_{\theta_{i}}(a)q(a)\right)\nabla_{\upsilon,\sigma}\log\mathbb{P}_{\upsilon,\sigma}(\theta_{i})\right]$ (41)
Simplifying and recognizing that $\sum_{a}\pi_{\theta_{i}}(a)q(a)=\mathbb{E}[R_{\tau}|\pi_{\theta_{i}}]$:
$=\sum_{\theta_{i}}\mathbb{P}_{\upsilon,\sigma}(\theta_{i})\left(\mathbb{E}[R_{\tau}|\pi_{\theta_{i}}]\right)\nabla_{\upsilon,\sigma}\log\mathbb{P}_{\upsilon,\sigma}(\theta_{i})$ (42)
Converting back to expectation form:
$=\mathbb{E}_{\theta\sim\rho}\left[\nabla_{\upsilon,\sigma}\log\mathbb{P}_{\upsilon,\sigma}(\theta)\cdot\mathbb{E}[R_{\tau}|\pi_{\theta}]\right]$ (43)
@@
Appendix D PB-SAC Algorithm
This appendix presents the complete algorithmic details of our PAC-Bayes Soft Actor-Critic (PB-SAC) method. Figure 3 provides a high-level illustration of the algorithm flow described in Section 4, showing the interaction between the actor, critic, and bound computation components. The periodic posterior update (step 2) guides exploration through PAC-Bayes bounds (step 3) while maintaining standard SAC training (step 1) using an adaptive sampling curriculum (step 4). Section D.1 provides the complete pseudocode with implementation details, including the posterior-guided exploration mechanism, the alternating bound optimization, and the integration with the critic adaptation loop.
@@ D.1 Pseudo Code
Init: Prior $\mu(\theta)=\mathcal{N}(\upsilon,\operatorname{diag}(\sigma^{2}))$
Init: $\text{actor\_frozen}=\text{False}$, prior moving average decay $\iota=0.99$
@@
6       $\theta_{\text{explore}}\leftarrow\arg\max_{\theta\sim\rho}Q(s,\pi_{\theta}(s))$
@@ 23    Update critics $Q_{1},Q_{2}$ using averaged targets
24
25    if adaptation steps completed then
26       $\text{actor\_frozen}\leftarrow\text{False}$
27
28      end if
29
30   end if
@@ 35    Compute discounted returns $G_{\text{IS}}$ with importance sampling on $\mathfrak{D}_{\text{train}}$
       /* Optimize posterior parameters for fixed $\kappa$ */
       $\sigma,\upsilon\leftarrow\arg\min_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\bm{(\sigma,\upsilon)}}}\mathcal{L}(\rho,\kappa)$
       // equation 9
       /* Optimize $\kappa$ for fixed posterior */
       $\kappa\leftarrow\arg\min_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\bm{\kappa^{\prime}}}}\mathcal{L}(\rho,\kappa^{\prime})$
       // equation 9
38
@@
    $\mu\leftarrow\iota\cdot\rho+(1-\iota)\cdot\mu$
@@
47    Linearly decay $\iota$
48
@@ 47    Linearly decay $\iota$
Appendix E More Results
This appendix presents extended experimental results that complement the main paper findings. Figure 4 shows detailed performance comparisons on the Walker2d-v5 environment, demonstrating both the algorithm comparisons (panel a) and the PAC-Bayes analysis (panel b). The empirical discounted return tracks closely with our certified lower bound, validating the theoretical guarantees in practice while consistently maintaining competitive performance in terms of rollout episodic return.
@@ Appendix E More Results
Appendix F Hyperparameter Selection
We carefully selected hyperparameters for our PAC-Bayes Soft Actor-Critic (PB-SAC) implementation to balance performance, sample efficiency, and theoretical guarantees. the common parameters with SAC are left unchanged, while we take the original hyperparameters of PBAC from the paper Tasdighi et al. (2025). Table 1 summarizes it all.

### v2

@@ Appendix D A note on the Markov assumption for Bellman errors
Appendix E REINFORCE Trick for Policy-Level Gradients
@@
Theorem E.1 (Policy-Level REINFORCE).
For a continuous posterior distribution $\rho_{\upsilon,\sigma}(\theta)=\mathcal{N}(\upsilon,\operatorname{diag}(\sigma^{2}))$ over policy parameters $\theta$, the gradient of the expected return can be computed as:
$\nabla_{\upsilon,\sigma}\mathbb{E}_{\theta\sim\rho}\left[\mathbb{E}[R_{t}|\pi_{\theta}]\right]=\mathbb{E}_{\theta\sim\rho}\left[\nabla_{\upsilon,\sigma}\log\rho_{\upsilon,\sigma}(\theta)\cdot\mathbb{E}[R_{t}|\pi_{\theta}]\right]$ (47)
@@
Let $J(\theta)=\mathbb{E}[R_{t}|\pi_{\theta}]$ denote the expected return of a specific policy parameterization $\theta$. We express the objective as an expectation over the continuous posterior density $\rho_{\upsilon,\sigma}(\theta)$ and apply the log-derivative trick directly:
$\nabla_{\upsilon,\sigma}\mathbb{E}_{\theta\sim\rho}\left[J(\theta)\right]$ $=\nabla_{\upsilon,\sigma}\int_{\Theta}\rho_{\upsilon,\sigma}(\theta)J(\theta)d\theta$ (48)
$=\int_{\Theta}\nabla_{\upsilon,\sigma}\rho_{\upsilon,\sigma}(\theta)J(\theta)d\theta$ (49)
$=\int_{\Theta}\rho_{\upsilon,\sigma}(\theta)\frac{\nabla_{\upsilon,\sigma}\rho_{\upsilon,\sigma}(\theta)}{\rho_{\upsilon,\sigma}(\theta)}J(\theta)d\theta\quad\text{(Log-derivative identity)}$ (50)
$=\int_{\Theta}\rho_{\upsilon,\sigma}(\theta)\nabla_{\upsilon,\sigma}\log\rho_{\upsilon,\sigma}(\theta)J(\theta)d\theta$ (51)
$=\mathbb{E}_{\theta\sim\rho}\left[\nabla_{\upsilon,\sigma}\log\rho_{\upsilon,\sigma}(\theta)\cdot J(\theta)\right]$ (52)
Substituting $J(\theta)$ back into the expression yields the theorem statement. ∎
@@
Appendix F PB-SAC Algorithm
This appendix presents the complete algorithmic details of our PAC-Bayes Soft Actor-Critic (PB-SAC) method. Figure 3 provides a high-level illustration of the algorithm flow described in Section 4, showing the interaction between the actor, critic, and bound computation components. The periodic posterior update (step 2) guides exploration through PAC-Bayes bounds (step 3) while maintaining standard SAC training (step 1) using an adaptive sampling (step 4). Section F.1 provides the complete pseudocode with implementation details, including the posterior-guided exploration mechanism, the alternating bound optimization, and the integration with the critic adaptation loop.
@@ F.1 Pseudo Code
Init: Prior $\mu(\theta)=\mathcal{N}(\upsilon_{0},\operatorname{diag}(\sigma_{0}^{2}))$
Init: $\text{actor\_frozen}=\text{False}$, prior moving average decay $\iota=0.99$, posterior sampling rate $\epsilon_{\text{explore}}=0.9$ (linearly decayed)
@@
6         Sample multiple policies $\{\theta_{i}\}\sim\rho$
7
8       $\theta_{\text{explore}}\leftarrow\arg\max_{\theta_{i}}Q(s,\pi_{\theta}(s))$
9
@@ 30  Update critics $Q_{1},Q_{2}$ using averaged targets
 If: adaptation steps completed then $\text{actor\_frozen}\leftarrow\text{False}$
32
@@ 38     Compute discounted returns $G_{\text{IS}}$ with importance sampling on $\mathfrak{D}_{\text{train}}$
       /* Optimize posterior parameters for fixed $\lambda$ */
       $\sigma,\upsilon\leftarrow\arg\min_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\bm{(\sigma,\upsilon)}}}\mathcal{J}(\rho,\lambda)$
       // inequality 14
41
       /* Optimize $\lambda$ for fixed posterior */
       $\lambda\leftarrow\arg\min_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\bm{\lambda^{\prime}}}}\mathcal{J}(\rho,\lambda^{\prime})$
       // inequality 14
42
@@
    $\upsilon_{0}\leftarrow\iota\cdot\upsilon+(1-\iota)\cdot\upsilon_{0};\quad\sigma_{0}\leftarrow\iota\cdot\sigma+(1-\iota)\cdot\sigma_{0}$
@@
51
52     Linearly decay $\iota$ and $\epsilon_{\text{explore}}$
@@
Appendix G More Results
This appendix presents extended experimental results that complement the main paper findings. Figure 4 shows detailed performance comparisons on Hopper-v5 and Walker2d-v5 environments, demonstrating both the algorithm comparisons (panel a) and the PAC-Bayes analysis (panel b). The empirical discounted return tracks closely with our certified lower bound, validating our analysis above.
@@ Appendix G More Results
G.1 Sparse-reward setting
(a) Algorithm comparisons
(b) PAC-Bayes analysis
(a) Mixing time: 1
(b) Mixing time: 50
(c) Mixing time: 100
(d) Mixing time: 300
(e) Mixing time: 500
(f) Mixing time: 1000
Figure 6: PAC-Bayesian bound behaviour with different fixed mixing time estimates on HalfCheetah-v5. Empirical discounted return (dashed) and certified discounted return (solid) are shown for mixing times ranging from 1 (a) to 1000 (f). Smaller mixing times yield tighter bounds but risk overconfidence if underestimated, while larger values provide conservative but still meaningful certificates. The algorithm maintains practical utility even with substantial mixing time overestimation, demonstrating robustness of the framework.
Appendix H Ablation study
(a) Adaptive Sampling
(b) Posterior-Guided Exploration
Figure 7: Ablations
Figure 8: KL divergence evolution
Appendix I Hyperparameter Selection
We carefully selected hyperparameters for our PAC-Bayes Soft Actor-Critic (PB-SAC) implementation to balance performance, sample efficiency, and theoretical guarantees. the common parameters with SAC are left unchanged, while we take the original hyperparameters of PBAC from the paper of Tasdighi et al. (2025). Table 1 summarizes it all.